# A Nuclear Interferometer for Ultra-Light Dark Matter Detection

### v1

@@ Abstract
We propose the nuclear interferometer - a single-photon interferometry experiment based upon the thorium-229 nuclear clock transition - as a novel detector for ultra-light dark matter (ULDM). Thanks to the enhanced sensitivity of this transition to the variation of fundamental constants, we find that possible realisations of such an experiment deploying either single ions or clouds of atoms have the potential to be complementary to advanced very-long-baseline terrestrial clock atom interferometers in the search for ultra-light dark matter with scalar couplings to photons. Nuclear interferometry may also offer an unparalleled window to new physics coupling to the QCD sector via quarks or gluons, with a discovery reach that could enhance existing and proposed experiments over a range of frequencies in the direction of well-motivated parameter space.
@@ I Introduction
ULDM could conceivably couple to the SM through a number of different ‘portals’, each of which has a distinct set of phenomenological consequences and in turn, experimental signatures. In this proposal we focus specifically on scalar ULDM with linear couplings to the SM. These couplings cause fundamental constants, and thus physical properties such as atomic energy levels and the distance scales which they govern, to oscillate in time at a frequency set by the DM mass, see e.g. Refs. [28, 29, 30, 31, 24]. This effect can be searched for in a variety of different ways including accelerometers [32], optical cavities [33], mechanical resonators [34, 35] and by comparing the frequency of two different atomic clocks [29, 31, 36, 37]. More recently, atom interferometers have been identified to offer the potential to probe ULDM-induced oscillations of atomic transition frequencies in a complementary region of parameter space to the aforementioned tests [38, 39]. In its most basic form, an atom interferometer constitutes an experiment which measures the phase difference between two spatially delocalised quantum superpositions of clouds of atoms [40]. Such experiments can be thought of analogously to optical interferometers but with the traditional beam-splitters and mirror elements realised via a sequence of controlled duration laser pulses which act to divide, reflect and recombine the atomic wave-packets by driving transitions between internal energy states. The presence of ULDM can be diagnosed by the existence of an oscillating phase difference between a pair of spatially separated single-photon [^1] Here, ‘single-photon’ refers specifically to a class of interferometers in which the desired transition is driven by a single laser frequency. This contrasts with interferometers based on Raman transitions which involve two distinct optical fields (see e.g. Ref. [41]). atom interferometers which are resonantly interrogated with common laser pulses [39]. Such multi-atom interferometer configurations are termed single-photon atom gradiometers and have the key advantage in that noise arising from fluctuations and jitters in the laser phase is common to both instruments and thus cancels in the differential measurement [42]. In order to achieve sensitivity to ULDM, such proposals look to both long ($\mathcal{O}$(km)) baselines and large-momentum-transfer (LMT) atom optics [^2] Such techniques (see e.g. Ref. [43]) effect an increase of the physical baseline by employing an increased number of laser pulses to augment the momentum transfer between the two arms of each atom interferometer and in turn the phase difference accumulated at the end of the sequence. [44] (see Sec. II.2) to increase the recorded phase difference. To sustain these practices it is essential for the pair of internal states on which the atom interferometer is based to be sufficiently long-lived such that spontaneous emission, which would otherwise exponentially degrade the sensitivity due to loss of coherence, is not significant. [^3] For a discussion of some of the potential implications of spontaneous decay on long-baseline atom interferometry in space, see Refs. [45, 46]. For the very long baselines and high number of sequential pulses demanded by ULDM searches (see e.g. [47, 48]), ultra-narrow transitions such as those deployed in optical lattice clocks are therefore typically taken advantage of. For this reason such atom interferometer realisations have earned the title of ‘large-momentum-transfer clock atom interferometry’ [49].
@@
In tandem, the discovery of an excited isomeric state in 229Th  [52, 53], has led to proposals for a ‘nuclear clock’ [54, 55, 56] based on transitions between the ground and the 229mTh excited isomer state. Arising due to accidental cancellations between contributions from strong and electromagnetic interactions [57, 58, 59], the small (8.3 eV [60]) gap in energy between the ground and excited states lies in the optical range [61] and offers an enhanced sensitivity to the variation of fundamental constants at the level of several orders of magnitude [57, 62, 63] over typical optical clock transitions in atoms, ions and highly charged ions [64]. Frequency comparisons between nuclear and atomic optical clocks could offer a highly sensitive probe of ULDM in the near future [65, 66, 67, 29, 68]. Due to rapid internal conversion, the excited isomeric state in electrically neutral 229Th is short-lived with a half-life of $\sim 7\,\mu$s [69]. Proposed realisations of the thorium nuclear clock for deployment in such experiments therefore look to either ionic forms [55] or to a thorium-doped lattice [70] in which internal conversion is forbidden and the isomeric state becomes long-lived with a radiative lifetime of $\sim 10^{4}$ s [65].
@@
We begin in Sec. II.1 by introducing the ULDM model that we will adopt as an example in this work, before computing the signal that it would imprint in a generic nuclear interferometry experiment in Sec. II.2. We consider potential noise sources in Sec. II.3 and use them to derive an expression for the reach of a general free-fall nuclear interferometer in terms of a handful of tunable parameters which characterise the experimental operating configurations. In Sec. III we explore two different ways, making use of single-ion clocks and clouds of atoms respectively, in which such a nuclear interferometer could be realised experimentally, and discuss how the aforementioned principles apply to each case. Forecasts of the sensitivity of both realisations to scalar ULDM with linear couplings to SM operators for a variety of experimental configurations that may become feasible in the future are presented in Sec. IV and compared to bounds from existing experiments and the predicted sensitivities of competing proposals. Some concluding remarks are offered in Sec. V.
@@ II.1 Ultra-light scalar Dark Matter
To motivate this proposal we consider an ultra-light singlet scalar field, $\phi$, which couples linearly [^4] Note that quadratic couplings also lead to the variation of fundamental constants [77], and become relevant if linear couplings are not present, forbidden for instance by a symmetry. We leave consideration of such couplings to future work. to the SM by way of non-derivative operators. We assume that this field comprises an $\mathcal{O}(1)$ fraction of the local DM abundance. Working in units in which $\hbar=c=1$ and following the notational conventions [^5] In this paramaterisation, the coefficients and normalisation of each of the terms in the Lagrangian are specifically chosen such that the dimensionless couplings are all RG-invariant [78]. of Refs. [78, 29] in parameterising the dimensionless couplings relative to Newton’s gravitational constant $G_{N}$, the leading-order low energy Lagrangian, applicable to scales $\gtrsim$ 1 GeV, can be expressed in terms of a relatively minimal set of operators as
@@
These effective operators feature in well-motivated models of ultra-light dark matter. For one example, in the presence of additional sources of CP violation beyond those included within the SM, both the QCD axion, and other pseudo-scalar axion-like particles possess CP-even scalar couplings in addition to their CP-odd pseudo-scalar interactions. In particular, the QCD axion can attain a non-zero scalar coupling to quarks [79, 80, 81], which is bounded from below by the CP violation in the SM, and from above by measurements of the neutron electric dipole moment.
@@ II.3 Experimental Sensitivity
Assuming that the sensitivity is limited by the quantum projection noise (or shot-noise) of the interrogated species [^7] Reaching shot-noise limited detection is a key design goal for atomic sensors and has already been demonstrated experimentally in atom interferometry - see e.g. Refs. [86, 87]. with spectral density $S_{\textnormal{n}}$, the minimum scalar coupling $d_{\phi}^{*}$ that could be detected at a signal-to-noise ratio SNR following a measurement period $T_{\textnormal{int}}$ is
@@ III Possible Realisations
We now discuss two distinct ways in which the principles outlined in the previous sections could be applied to perform interferometry on the ground-excited isomeric state transition in 229Th - the first involving single-ion ‘clocks’, and the second clouds of neutral atoms. Both scenarios present a distinct set of experimental challenges from the sources traditionally considered for long-baseline interferometry and yet, as will be shown, may offer the potential to open up new phenomenological territory. Differing both in their physical nature and the properties of the ground-excited state transition, these two realisations demand starkly different operating configurations (as characterised by the parameters identified in Sec. II.3) which we now outline.
@@ III.1 Single Ions
As previously discussed, conventional ULDM searches using atom interferometry propose to make use of ultra-narrow clock transitions which, thanks to long excited state lifetimes, allow for the use of long baselines and large inter-arm momentum transfers. Although the excited isomeric state lifetime for neutral 229Th is only $\sim 10^{-5}$ s due to internal conversion, when ionised, this decay channel becomes inaccessible and the (now radiative decay limited) lifetime is of order $\sim 10^{4}$ s, falling within the remits of clock interferometry. We thus begin our exploration of the potential usage of 229Th in free-fall matter-wave interferometry by considering working with an ionic [^8] From the perspective of the interferometry sequence, the precise charge of the ion is unimportant and we thus leave this unspecified. form.
@@
In addition to deploying a high number of atoms per shot, the sensitivity forecasts for long-baseline atom interferometers also make use of multiplexing (see e.g. Ref. [49]). Exploiting the Doppler shift to the transition frequency experienced by clouds of atoms moving at different velocities, this technique allows the time interval between successive shot launches to be reduced below the interferometer period (2T) with multiple independent interferometers concurrently in operation. Since (in typical Mach-Zehnder interferometry schemes) the atoms are launched upwards and subsequently fall under gravity, the trajectories of atoms in successive clouds cross. The atoms, which can be treated ballistically, are effectively non-interacting such that this trajectory overlap can occur without disrupting the generation of interference at detection. The same can not be said for ions due to electrostatic repulsion. This imposes a lower limit on the time interval between shots: $\Delta t\geq 2T$. We note that whilst this condition applies to interferometry configurations in which the ions are launched upwards and latterly fall under gravity as in this proposal, if instead a configuration could be utilised in which they only travelled in one direction (i.e. if dropped from rest with the primary laser pulses fired vertically downwards) such that they were detected in a different region from launch, it could be possible to achieve a higher effective shot flux due to the possibility of now applying multiplexing. In this case the experimental parameters need to be carefully selected in order to ensure that there are no intersections between the trajectories of each of the arms of the two interferometers. In order for such linear, non-intersecting trajectories to simultaneously satisfy the geometric constraints of the baseline, smaller $T$, $n$ and $\Delta z$ than the conventional set-up are likely to be necessary, somewhat suppressing the potential sensitivity gain associated with a higher ion flux. We leave a more detailed exploration of the potential benefit of such scenarios to future work.
@@
For the terrestrial setup, we therefore consider a 1 km baseline with $n$ = 2500 LMT optics. We set the interferometer separation $\Delta z$, to its maximum value
@@
We model our space-based instrument on the design considered in Refs. [92, 42, 47, 93, 94] comprising two satellites in medium Earth orbit separated by a distance of 4.4 ×$\times 10^{7}$ m along a single line-of-sight. In design A we take the more conservative assumption that the region over which the ions are interrogated is confined to the interior of the satellites. This constrains the maximum wavepacket separation during the interferometric sequence, $v_{n}T$, to be less than $\sim$ 1 m [47, 92]. In contrast, if the interrogation region could be located outside the satellite (as in the designs of Ref. [95]), wave-packet separations on the order of $\sim 100$ m could be achieved. Whilst these inequalities are linear in both $T$ and $n$, we note that due to the dependence of the ion flux and thus shot-noise spectral density on $T$, it is preferable to deploy larger $n$. We select $n$ and $T$ to satisfy these inequalities, accounting for the fact that $n$ should be even and that the total duration of the interferometer sequence should exceed the total time taken for the (sequentially fired) laser pulses to traverse the baseline, i.e. $2T>(4n-1)L/c$ , where $c$ is the speed of light. Whilst the parameters used in our forecasts are representative of the reach that could be attained in such experiments we note that further optimisation could be performed if one had a particular signal or region of parameter space in mind.
For each of the three designs we set $S_{\textnormal{n}}$ to be consistent with Eq. 19, accounting for the fact that one must have both $N_{\textnormal{s}}=1$ and $\Delta t\geq 2T$.
@@
Terrestrial 1000 1.2 2500 2.3 965
Space-based (A) $4.4\times 10^{7}$ 5 16 4.7 $4.4\times 10^{7}$
Space-based (B) $4.4\times 10^{7}$ 70 120 17.6 $4.4\times 10^{7}$
@@
Table 1: The experimental parameters characterising the single-ion nuclear interferometers considered in this work. In the space-based configuration A the interrogation region is confined to be within the satellites, whereas in configuration B it is exterior.

### v2

@@ Abstract
We propose the nuclear interferometer - a single-photon interferometry experiment based upon the thorium-229 nuclear clock transition - as a novel detector for ultra-light dark matter (ULDM). Thanks to the enhanced sensitivity of this transition to the variation of fundamental constants, we find that possible realisations of such an experiment deploying either single ions or clouds of atoms have the potential to complement advanced very-long-baseline terrestrial clock atom interferometers in the search for ultra-light dark matter with scalar couplings to photons in the future. Nuclear interferometry may also offer an unparalleled window to new physics coupling to the QCD sector via quarks or gluons, with a discovery reach that could enhance existing and proposed experiments over a range of frequencies in the direction of well-motivated parameter space.
@@ I Introduction
ULDM could conceivably couple to the SM through a number of different ‘portals’, each of which has a distinct set of phenomenological consequences and in turn, experimental signatures. In this proposal we focus specifically on scalar ULDM with linear couplings to the SM. These couplings cause fundamental constants, and thus physical properties such as atomic energy levels and the distance scales which they govern, to oscillate in time at a frequency set by the DM mass, see e.g. Refs. [28, 29, 30, 31, 24]. This effect can be searched for in a variety of different ways including accelerometers [32], optical cavities [33], mechanical resonators [34, 35] and by comparing the frequency of two different atomic (see e.g. Refs. [29, 31, 36, 37]) or molecular [38] clocks. More recently, atom interferometers have been identified to offer the potential to probe ULDM-induced oscillations of atomic transition frequencies in a complementary region of parameter space to the aforementioned tests [39, 40]. In its most basic form, an atom interferometer constitutes an experiment which measures the phase difference between two spatially delocalised quantum superpositions of clouds of atoms [41]. Such experiments can be thought of analogously to optical interferometers but with the traditional beam-splitters and mirror elements realised via a sequence of controlled duration laser pulses which act to divide, reflect and recombine the atomic wave-packets by driving transitions between internal energy states. The presence of ULDM can be diagnosed by the existence of an oscillating phase difference between a pair of spatially separated single-photon [^1] Here, ‘single-photon’ refers specifically to a class of interferometers in which the desired transition is driven by a single laser frequency. This contrasts with interferometers based on Raman transitions which involve two distinct optical fields (see e.g. Ref. [42]). atom interferometers which are resonantly interrogated with common laser pulses [40]. Such multi-atom interferometer configurations are termed single-photon atom gradiometers and have the key advantage in that noise arising from fluctuations and jitters in the laser phase is common to both instruments and thus cancels in the differential measurement [43]. In order to achieve sensitivity to ULDM, such proposals look to both long ($\mathcal{O}$(km)) baselines and large-momentum-transfer (LMT) atom optics [^2] Such techniques (see e.g. Refs. [44, 45, 46]) effect an increase of the physical baseline by employing an increased number of laser pulses to augment the momentum transfer between the two arms of each atom interferometer and in turn the phase difference accumulated at the end of the sequence. [47] (see Sec. II.2) to increase the recorded phase difference. To sustain these practices it is essential for the pair of internal states on which the atom interferometer is based to be sufficiently long-lived such that spontaneous emission, which would otherwise exponentially degrade the sensitivity due to loss of coherence, is not significant. [^3] For a discussion of some of the potential implications of spontaneous decay on long-baseline atom interferometry in space, see Refs. [48, 49]. For the very long baselines and high number of sequential pulses demanded by ULDM searches (see e.g. [50, 51]), ultra-narrow transitions such as those deployed in optical lattice clocks are therefore typically taken advantage of. For this reason such atom interferometer realisations have earned the title of ‘large-momentum-transfer clock atom interferometry’ [52].
@@
In tandem, the discovery of an excited isomeric state in 229Th  [56, 57], has led to proposals for a ‘nuclear clock’ [58, 59, 60] based on transitions between the ground and the 229mTh excited isomer state. Arising due to accidental cancellations between contributions from strong and electromagnetic interactions [61, 62, 63], the small (8.3 eV [64]) gap in energy between the ground and excited states lies in the optical range [65] and offers an enhanced sensitivity to the variation of fundamental constants at the level of several orders of magnitude [61, 66, 67, 68] over typical optical clock transitions in atoms, ions and highly charged ions [69]. Frequency comparisons between nuclear and atomic optical clocks could offer a highly sensitive probe of ULDM in the near future [70, 71, 72, 29, 73], and the (still broad) lineshape already constrains the ULDM parameter space [74]. Due to rapid internal conversion, the excited isomeric state in electrically neutral 229Th is short-lived with a half-life of $\sim 7\,\mu$s [75]. Proposed realisations of the thorium nuclear clock for deployment in such experiments therefore look to either ionic forms [59] or to a thorium-doped lattice [76] in which internal conversion is forbidden and the isomeric state becomes long-lived with a radiative lifetime of $\sim 10^{4}$ s [70].
@@
We begin in Sec. II.1 by introducing the ULDM model that we will adopt as an example in this work, before computing the signal that it would imprint in a generic nuclear interferometry experiment in Sec. II.2. In Sec. II.3 we then derive an expression for the reach of a general free-fall nuclear interferometer assuming shot-noise limited detection, in terms of a handful of tunable parameters which characterise the experimental operating configurations. In Sec. III we explore two different ways, making use of single-ion clocks and clouds of atoms respectively, in which such a nuclear interferometer could be realised experimentally and discuss how the aforementioned principles apply to each case. The challenges of working with both systems are discussed and any requisite technological advancements outlined. Forecasts of the sensitivity of both realisations to scalar ULDM with linear couplings to SM operators for a variety of experimental configurations that may become feasible in the future are presented in Sec. IV and compared to bounds from existing experiments and the predicted sensitivities of competing proposals. Some concluding remarks are offered in Sec. V.
@@ II.1 Ultra-light scalar Dark Matter
To motivate this proposal we consider an ultra-light singlet scalar field, $\phi$, which couples linearly [^4] Note that quadratic couplings also lead to the variation of fundamental constants [83], and become relevant if linear couplings are not present, forbidden for instance by a symmetry. We leave consideration of such couplings to future work. to the SM by way of non-derivative operators. We assume that this field comprises an $\mathcal{O}(1)$ fraction [^5] An exploration of the ability of scalar couplings to generate the correct relic abundance can be found in Ref. [84]. of the local DM abundance. Working in units in which $\hbar=c=1$ and following the notational conventions [^6] In this paramaterisation, the coefficients and normalisation of each of the terms in the Lagrangian are specifically chosen such that the dimensionless couplings are all RG-invariant [85]. of Refs. [85, 29] in parameterising the dimensionless couplings relative to Newton’s gravitational constant $G_{N}$, the leading-order low energy Lagrangian, applicable to scales $\gtrsim$ 1 GeV, can be expressed in terms of a relatively minimal set of operators as
@@
These effective operators feature in several well-motivated models of ultra-light dark matter. For one example, in the presence of additional sources of CP violation beyond those included within the SM, both the QCD axion, and other pseudo-scalar axion-like particles possess CP-even scalar couplings in addition to their CP-odd pseudo-scalar interactions. In particular, the QCD axion can attain a non-zero scalar coupling to quarks [86, 87, 88], which is bounded from below by the CP violation in the SM, and from above by measurements of the neutron electric dipole moment.
@@ II.3 Experimental Sensitivity
Assuming that the sensitivity is limited by the quantum projection noise (or shot-noise) of the interrogated species with spectral density $S_{\textnormal{n}}$, the minimum scalar coupling $d_{\phi}^{*}$ that could be detected at a signal-to-noise ratio SNR following a measurement period $T_{\textnormal{int}}$ is
@@
Before moving on we comment on the assumption that the experimental sensitivity is shot-noise limited. Achieving shot-noise limited detection is a key design goal for any atomic sensor and requires, by definition, that all other sources of noise arising from both systematic and stochastic effects, are subdominant. Whilst this ideal has already been demonstrated experimentally for small-scale table-top atom interferometers (see e.g. Refs. [96, 97, 98]), scaling the experiments up to longer baselines brings additional considerations and challenges for noise control. Comprehensive assessments of the many sources of noise relevant to long-baseline atom interferometry (e.g. gravity gradients, laser phase noise, mechanical vibrations etc.), in addition to the various practices and techniques that may be deployed to control these, have been discussed at length in literature (see e.g. Refs. [99, 100, 54, 101]) and used to place requirements on various aspects of the experimental operation. Where not already achievable with current technologies, these stand as targets for ongoing research and development. Given that the arguments for many of these noise sources carry over straightforwardly to both of the nuclear interferometry realisations presented in this work, we do not attempt to replicate this discussion, assuming that such conditions and practices could also be successfully applied to any future nuclear interferometer experiment. Instead, in Sec. III we identify just the most significant sources of noise that are likely to impact each of the proposals outlined in this work, compare them to the shot-noise floor and discuss any resulting implications for the instrument design or experimental operation.
@@ III Possible Realisations
We now discuss two distinct ways in which the principles outlined in the previous sections could be applied to perform interferometry on the ground-excited isomeric state transition in 229Th in the future - the first involving single-ion ‘clocks’, and the second clouds of neutral atoms. Both scenarios present a distinct set of experimental challenges from the sources traditionally considered for long-baseline interferometry, yet, as will be shown, may offer the potential to open up new phenomenological territory in the future provided the requisite technologies can be developed. Differing both in their physical nature and the properties of the ground-excited state transition, these two realisations demand starkly different operating configurations (as characterised by the parameters identified in Sec. II.3) which we now outline.
@@ III.1 Single Ions
As previously discussed, conventional ULDM searches using atom interferometry propose to make use of ultra-narrow clock transitions which, thanks to long excited state lifetimes, allow for the use of long baselines and large inter-arm momentum transfers. Although the excited isomeric state lifetime for neutral 229Th is only $\sim 10^{-5}$ s [75] due to internal conversion, when ionised, this decay channel becomes inaccessible and the (now radiative decay limited) lifetime is of order $\sim 10^{4}$ s [77, 102], falling within the remits of clock interferometry. We thus begin our exploration of the potential usage of 229Th in free-fall matter-wave interferometry by considering working with an ionic [^8] From the perspective of the interferometry sequence, the precise charge of the ion is unimportant and we thus leave this unspecified. form.
@@
In addition to deploying a high number of atoms per shot, the sensitivity forecasts for long-baseline atom interferometers also make use of multiplexing (see e.g. Ref. [52]). Exploiting the Doppler shift to the transition frequency experienced by clouds of atoms moving at different velocities, this technique allows the time interval between successive shot launches to be reduced below the interferometer period (2T) with multiple independent interferometers concurrently in operation. Since (in typical Mach-Zehnder interferometry schemes) the atoms are launched upwards and subsequently fall under gravity, the trajectories of atoms in successive clouds cross. The atoms, which can be treated ballistically, are effectively non-interacting such that this trajectory overlap can occur without disrupting the generation of interference at detection. The same can not be said for ions due to electrostatic repulsion. This imposes a lower limit on the time interval between shots: $\Delta t\geq 2T$. We note that whilst this condition applies to interferometry configurations in which the ions are launched upwards and latterly fall under gravity as in this proposal, if instead a configuration could be utilised in which they only travelled in one direction (i.e. if dropped from rest with the primary laser pulses fired vertically downwards) such that they were detected in a different region from launch, it could be possible to achieve a higher effective shot flux due to the possibility of now applying multiplexing. In this case the experimental parameters would need to be carefully selected in order to ensure that there are no intersections between the trajectories of each of the arms of the two interferometers. In order for such linear, non-intersecting trajectories to simultaneously satisfy the geometric constraints of the baseline, smaller $T$, $n$ and $\Delta z$ than the conventional set-up are likely to be necessary, somewhat suppressing the potential sensitivity gain associated with a higher ion flux. We leave a more detailed exploration of the potential benefit of such scenarios to future work.
@@
For the terrestrial setup, we therefore consider a 1 km baseline with $n$ = 1000 LMT optics. We set the interferometer separation $\Delta z$, to its maximum value
@@
We model our space-based instrument on the design considered in Refs. [104, 43, 50, 105, 106], comprising two satellites in medium Earth orbit separated by a distance of 4.4×$\times 10^{7}$ m along a single line-of-sight. We take the conservative assumption that the region over which the ions are interrogated is confined to within the interior of the satellites. This constrains the maximum wavepacket separation during the interferometric sequence, $v_{n}T$, to be less than $\sim$ 1 m [50, 104]. Whilst this inequality is linear in both $T$ and $n$, we note that due to the dependence of the ion flux and thus the shot-noise spectral density on $T$, it is preferable to deploy larger $n$. We select $n$ and $T$ to satisfy this inequality, accounting for the fact that $n$ should be even and that the total duration of the interferometer sequence should exceed the total time taken for the (sequentially fired) laser pulses to traverse the baseline, i.e. $2T>(4n-1)L/c$, where $c$ is the speed of light. Whilst the parameters used in our forecasts are representative of the reach that could be attained in such an experiment we note that further optimisation could be performed if one had a particular signal or region of parameter space in mind. Although more ambitious space-based designs in which the interrogation takes place exterior to the satellite allowing for wavepacket separations on the order of $\sim 100$ m have been proposed [99], this removes the possibility of applying magnetic shielding and would thus leave the ions subject to uncontrollable electric and magnetic fields. This would likely be a major source of gradiometer phase noise which is exacerbated by the large interrogation times associated with this set-up [^9] As will be discussed in due course, the gradiometer phase shift induced by electromagnetic field noise scales with $T^{2}$. and we therefore deem such a design unsuitable and do not consider it further.
For both the terrestrial and space-based designs we set $S_{\textnormal{n}}$ to be consistent with Eq. 19, accounting for the fact that one must have both $N_{\textnormal{s}}=1$ ions per shot and $\Delta t\geq 2T$.
@@
Terrestrial 1000 1.2 1000 2.3 986
Space-based $4.4\times 10^{7}$ 5 16 4.7 $4.4\times 10^{7}$
@@
Table 1: The experimental parameters characterising the single-ion nuclear interferometers considered in this work: the baseline $L$, the interrogation time $T$, the LMT order $n$, the shot-noise spectral density $S_{\textnormal{n}}$, and the vertical separation $\Delta z$ between the two interferometers. In the space-based configuration, the interrogation region is taken to be within the interior of the satellites.
With ions it is also necessary to factor in their enhanced sensitivity to electromagnetic fields. Conventional atom interferometers already have stringent requirements on their internal magnetic environments in order to suppress unwanted contributions to the gradiometer phase which arise from (quadratic) Zeeman shifts to the transition frequency in the presence of temporally or spatially varying magnetic fields. For these interferometers there are two main sources of time-dependent magnetic fields that need to be considered: fluctuations in the local magnetic environment (i.e. for terrestrial interferometers the Earth’s magnetic field), and variations in the bias magnetic field which is applied to set a quantisation axis for the atoms and enable the use of magnetically insensitive states for which first order Zeeman shifts are not present. Whilst the use of a bias field is not necessary for nuclear interferometry, [^10] As will be discussed, the nuclear transition frequency is predicted to be largely insensitive to perturbations from external fields such that Zeeman shifts are strongly suppressed. fluctuations in the environmental magnetic field remain a potential concern.
The transition frequency of proposed realisations of ionic nuclear clocks is predicted to be significantly more stable to perturbations from external fields than typical atomic optical clocks [59]. Instead, the dominant effect of an external magnetic field on the gradiometer phase arises due to perturbations to the ion’s trajectory. For a single interferometer, the leading order phase difference, $\Phi_{i}$, comes from the acceleration of the interrogated species in the direction of the laser wavevector (i.e. the $z$ direction ): $\Phi_{i}\sim n\omega_{N}a_{z}T^{2}/c$, where $a_{z}$ denotes the $z$-component of the acceleration experienced by the ions and $k$ is the magnitude of the laser wavevector [107, 100, 43]. A magnetic field of magnitude $B$ orientated perpendicularly to the $z$-axis induces an additional vertical acceleration of $a_{z}=qBv_{t}/m_{\textnormal{th}}$ on an ion of charge $q$ and mass $m_{\textnormal{th}}$, where $v_{t}$ denotes the speed of the particle in the transverse direction perpendicular to the magnetic field. Whilst, in the case of spatially and temporally uniform fields, the phase accrued by each isolated interferometer due to this effect is identical and thus cancels in a differential measurement, spatially dependent magnetic field fluctuations would produce spurious gradiometer phase noise, which, if located in the instrument frequency band, could be indistinguishable from an oscillating ULDM signal.
In particular, if the two interferometers experience a difference in magnetic field of $\delta B$ in some direction perpendicular to the $z$ axis over the course of the interferometer sequence, an additional gradiometer phase shift of
$\begin{split}\delta\Phi\sim\left(1\,\textnormal{Hz}^{-\frac{1}{2}}\right)\left(\frac{q}{e}\right)\left(\frac{n}{1000}\right)\left(\frac{v_{t}}{0.1\textnormal{mm/s}}\right)\times\\ \left(\frac{\delta B}{\textnormal{pT}/\sqrt{\textnormal{Hz}}}\right)\left(\frac{T}{\textnormal{s}}\right)^{2}~{},\end{split}$ (22)
where $e$ = $1.6\times 10^{-19}$ C is the charge of the electron, will be induced. Assuming the transverse velocities of the ions to be $\sim 0.1$ mm/s corresponding to an effective temperature [^11] Whilst techniques for cooling single trapped ions down to $\sim$ nK thermal temperatures have been proposed [108], we note that transverse velocities of $\sim 0.1$ mm/s could be realised at higher thermal temperatures by allowing for a greater longitudinal velocity. It is the use of such cooling practices in conventional atom interferometry which have given rise to the use of the notion of ‘effective’ temperature - the thermal temperature corresponding to the speed in the transverse plane. on the order of $\sim$ nK, the terrestrial and space-based configurations detailed in Tab. 1 remain shot-noise limited for magnetic field fluctuations of amplitude $\delta B\leq 1$ $\textnormal{pT}/\sqrt{\textnormal{Hz}}$ and $\delta B\leq 7$ $\textnormal{pT}/\sqrt{\textnormal{Hz}}$, respectively. This expression assumes that the time scale over which the field fluctuates is long relative to the interrogation time such that the magnetic field experienced by each interferometer remains approximately constant over the interferometer sequence, which is reasonable for the sub-Hz frequencies and $\mathcal{O}$(s) interrogation times of relevance here.
It remains to compare these requirements to the magnetic field backgrounds that may be expected for these experiments. At present, the techniques for shielding magnetic fields over the large scales necessary for terrestrial very-long-baseline interferometry (i.e. 10’s - 100’s of metres) offer screening down to the $\sim$ nT level [109, 110]. [^12] Whilst the damping of magnetic field fluctuations and non-uniformities to sub-pT levels has been achieved in certain laboratory settings (see e.g. Ref. [111]), the shielding required for very-long-baseline interferometry is significantly more challenging owing both to the total shield length required [54] and to problems of flux leakage in large length-to-diameter ratio shields [112]. The amplitudes of stochastic fluctuations in the terrestrial magnetic field at the surface of the Earth in the frequency band of interest are expected to be below these levels [113, 114] and would therefore evade present technologies. Whilst at frequencies exceeding $\sim$ 1 Hz, the magnitude of these fluctuations is sufficiently small to satisfy the condition for shot-noise domination, they rapidly increase in magnitude below this point [113, 114] such that in the absence of any other means of mitigation, the experiment would become magnetic noise limited with current technology. For the experiment to remain shot-noise limited across the entire operational frequency band, a three order of magnitude improvement in the shielding factor of large scale magnetic shielding would be required. Even without such advances however it may be possible to substantially reduce the impact of magnetic noise through active surveillance and mitigation of the local magnetic environment. To this end, it is possible to measure magnetic fields to sub pT levels with magnetometers. Using an array of such instruments to monitor the local magnetic field strength and direction may allow for a reconstruction of the ion trajectories and the subsequent computation and subtraction of the corresponding gradiometer phase response, as outlined in Ref. [99].
In contrast, temporal fluctuations in the interplanetary magnetic field (as relevant to a space-based instrument) of frequency $f$ at $\sim$1 AU have been measured to be $\sim 0.1\frac{\textnormal{nT}}{\sqrt{\textnormal{Hz}}}\times\left(\frac{10^{-2}\textnormal{Hz}}{f}\right)$ [115]. If these fluctuations could not be shielded or damped, the proposed space-based nuclear interferometer would become limited by magnetic field noise at frequencies below $\sim 0.1$ Hz. To this end, retaining shot-noise limited detection over the entire frequency band of the experiment demands either shielding capable of suppressing fluctuations down to the $\sim$ 7 pT level across the $\sim$ 1 m interrogation region, or a successful active monitoring and mitigation campaign [99]. Given the substantially smaller scales over which fluctuations need to be screened, meeting these shielding demands is likely to prove less challenging than those of the terrestrially-based case.
Gradiometer phase noise can similarly arise from spurious vertical accelerations induced by temporally fluctuating electric fields. Specifically, a difference in the $z$-component of the electric field experienced by each interferometer of $\delta E$ generates a gradiometer phase shift of
$\delta\Phi\sim\left(1\,\textnormal{Hz}^{-\frac{1}{2}}\right)\left(\frac{q}{e}\right)\left(\frac{n}{1000}\right)\left(\frac{\delta E}{\textnormal{0.1\,fVm}^{-1}/\sqrt{\textnormal{Hz}}}\right)\left(\frac{T}{\textnormal{s}}\right)^{2}~{}.$ (23)
[…]
Local patch potentials, spatial and temporal variations of electrostatic potentials across surfaces [116], which are responsible for short-range forces close to surfaces, may pose a further challenge. Given that the dominant way in which electric fields are imprinted in the gradiometer phase comes from perturbations to the $z$-component of the acceleration of the ions, it is vertical E-fields, as generated from horizontal surfaces which are of greatest concern. In this respect, the ion trajectories (as determined by their launch heights and velocities) should be chosen to keep the ions sufficiently far from the surfaces at the upper and lower ends of the chamber.